OpenStudy (anonymous):

Mixture Problem: Addison is mixing hummingbird food to fill her feeders. Ideally, the hummingbird food should contain 25% sugar. Addison remembers that she has 8 ounces of leftover hummingbird food that contains 35% sugar. She decides to mix some diluted food to add to the leftover food. If the diluted food contains 10% sugar, how much should she add to the food with 35% sugar to get the ideal hummingbird food with 25% sugar? a. What equation is used to solve this problem? b. Identify the given information. How can you represent each food mixture with an expression? What amount will you repr

OpenStudy (anonymous):

yay someones looking at my question

OpenStudy (anonymous):

use 100*(ounces of sugar)/(ounces of bird feed) = (percentage of sugar) If the 8 ounces of bird feed with 35 percent sugar, then it must have 2.8 ounces of sugar.

OpenStudy (anonymous):

is that the answer??

OpenStudy (anonymous):

Now solve (2.8+.1*x)/(8+x)=25/100

OpenStudy (anonymous):

o solve a problem involving the sum of two mixtures use the equation, (amount mix A)(% mix A) + (amount mix B) (% mix B) = (amount resulting mix)(% resulting mix). Enter the given information into the equation and solve for the unknown variable

OpenStudy (anonymous):

my lesson says to do that is that what u did

OpenStudy (anonymous):

.1 represents the 10 percents diluted sugar

OpenStudy (anonymous):

and x is the answer?

OpenStudy (anonymous):

x is the amount of diluted birdfeed she must add. 5.333 ounces.

OpenStudy (anonymous):

so the equation you had was right.

OpenStudy (anonymous):

i hope that helps

OpenStudy (anonymous):

alright thanks so much

OpenStudy (anonymous):

idk if this answers right so help u guys i need second opinions

OpenStudy (mathteacher1729):

This page has lots of worked examples and it gives a nice table which might help organize your work: http://www.onlinemathlearning.com/mixture-problems.html (click example 1, 2, or 3)

OpenStudy (anonymous):

thanks do you think the answer the person gave is right though

OpenStudy (mathteacher1729):

Fill in the table: \[\begin{array}{r|ccc}sugar & original & added & result \\ \hline \\ concentration (\%) & & \\amount (oz.)\end{array}\]

OpenStudy (anonymous):

I'm asking if the answers right i don't care how u work it out

OpenStudy (mathteacher1729):

Your problem is almost identical to this one: http://www.onlinemathlearning.com/mixture-problems.html#add Only you are using SUGAR instead of SALT. :)

hero (hero):

Maybe a visual diagram would work: